# Deep Learning-Based Breast Cancer Detection in Mammography: A Multi-Center Validation Study in Thai Population

### v1

@@ 3.4 Model Details
Our deep learning model consists of three parts: an encoder, a decoder, and aggregation layers. The encoder network is based on the EfficientNetV2(S) [11] architecture, the decoder module uses upsampling layers to increase the size of input feature maps from the encoder, and the aggregation layer applies Probabilistic-Class Activation Map or PCAM [14] pooling to combine final feature maps to classification and heatmap outputs.
@@
A confidence score and heatmap for each class is estimated by the model. Suspicion score per class (Calcification, Mass, Other) by breast (left/right) is presented to the end users. Heatmaps suggesting the location of suspicious lesions are also overlaid on the final image. Heatmaps are also generated for benign lesions and are visually represented using distinct styles to subtly alert users to their presence. In color heatmaps, benign lesions are depicted using a blue color scheme, while in greyscale heatmaps, they are represented by dotted lines. These visual cues aim to draw the user’s attention to the benign lesions without overemphasizing their importance.
@@ 4 Clinical Evaluation
The in-domain test set is a part of the Siriraj dataset kept aside for evaluation of the model. Pathologically or biopsy-confirmed dataset includes mammogram-guided biopsy cases and two-year follow-up confirmed benign and normal samples. For assessing out-of-domain generalizability, we used a diverse dataset collected from various hospitals. Table 3 provides detailed information about these datasets.
@@ 4.3 Evaluation on out-of-domain dataset
With an AUC of 0.937, our model demonstrates excellent performance on this dataset as well. The overall sensitivity and specificity of 93.0% and 68.7% for cancer detection is a bit decreased but still in line with other datasets.
@@ 5.1 Concordance Test
For localization concordance, heatmaps from the MMG model were compared with the localization ground truths provided by the radiologists and then divided into one of the four categories based on the agreements between them, as shown in Table 7. The MMG model heatmaps were presented to the radiologists as a complete representation of the localization of all lesions, without considering the specific lesion types. While previous work [[17]] focused on classification concordance, we have extended this approach to include localization results as well.
@@ 6 Conclusion
Availability of data and materials
The datasets in the current study are not publicly available due to patient data privacy concerns, but are available from the corresponding author on reasonable request.
Competing interests
The authors declare that they have no personal or professional conflicts of interest regarding any aspect of this study.

### v2

@@ 3.4 Model Details
Figure 2 illustrates the architecture of our deep learning model. Our deep learning model consists of three parts: an encoder, a decoder, and aggregation layers. The encoder network is based on the EfficientNetV2(S) [11] architecture, the decoder module uses upsampling layers to increase the size of input feature maps from the encoder, and the aggregation layer applies Probabilistic-Class Activation Map or PCAM [14] pooling to combine final feature maps to classification and heatmap outputs.
@@
A confidence score and heatmap for each class is estimated by the model. Suspicion score per class (Calcification, Mass, Other) by breast (left/right) is presented to the end users. Heatmaps suggesting the location of suspicious lesions are also overlaid on the final image. Heatmaps are also generated for benign lesions and are visually represented using distinct styles to subtly alert users to their presence. In color heatmaps, benign lesions are depicted using a blue color scheme, while in greyscale heatmaps, they are represented by dotted lines. These visual cues aim to draw the user’s attention to the benign lesions without overemphasizing their importance. Figure 3 shows examples of the final output images with confidence scores and overlaid heatmaps.
@@ 4 Clinical Evaluation
The in-domain test set is a part of the Siriraj dataset kept aside for evaluation of the model. Pathologically or biopsy-confirmed dataset includes mammogram-guided biopsy cases and two-year follow-up confirmed benign and normal samples. For assessing out-of-domain generalizability, we used a diverse dataset collected from various hospitals. Detailed information about these datasets is provided in Table 3 and Table 4.
@@ 4.3 Evaluation on out-of-domain dataset
Table 9 and Table 10 shows the performance on the generalization set. With an AUC of 0.937, our model demonstrates excellent performance on this dataset as well. The overall sensitivity and specificity of 93.0% and 68.7% for cancer detection shows a slight decrease but remains clinically acceptable, indicating robust generalizability across different clinical environments.
The model’s localization performance on out-of-domain data (Table 10) shows varying results across lesion types. While calcification detection maintained strong performance (LLF 0.788), mass localization showed some degradation (LLF 0.657), suggesting that mass appearance may vary more significantly across different hospital settings due to differences in imaging protocols or patient populations. The relatively small decrease in overall performance metrics compared to the biopsy-confirmed dataset demonstrates the model’s ability to maintain clinical utility when deployed in new healthcare settings, though initial monitoring of performance during deployment would be advisable.
@@ 5.1 Concordance Test
For localization concordance, heatmaps from the MMG model were compared with the localization ground truths provided by the radiologists and then divided into one of the four categories based on the agreements between them, as shown in Table 7. The MMG model heatmaps were presented to the radiologists as a complete representation of the localization of all lesions, without considering the specific lesion types. While previous work [17] focused on classification concordance, we have extended this approach to include localization results as well. The detailed criteria for each concordance type are shown in Table 11 and Table 12.